OpenStudy (anonymous):

Evaluate the integral (tangent inverese of x/(x)^2)dx

OpenStudy (anonymous):

Well if the question is integral of \[\tan^{-1} (x/x^{2})\] Then that equals to integral of \[\tan^{-1} (1/x)\] which by using the formula for integral of \[\tan^{-1} (x)\] =\[1/\sqrt{1+x^2}\] here it will be \[1/\sqrt{1+(1/x^2)} = x/\sqrt{x^2+1}\]

OpenStudy (anonymous):

no is tan^-1(x) over x^2

OpenStudy (anonymous):

you dont know it ?

OpenStudy (anonymous):

integral of [tan^-1(x)]/x^2 dx =integral of (x^-2)[tan^-1(x)] dx = -(x^-1)[tan^-1(x)] +ln[(x)/(1+x^2)]+C

OpenStudy (anonymous):

simagholami do you have answer choices?